OpenStudy (anonymous):

What is the probability of obtaining exactly seven heads in eight flips of a coin, given that at least one is a head?

OpenStudy (anonymous):

1/2 x 1/2 x 1/2 x 1/2 x 1/2 x 1/2 x 1/2 = 1/2^7 so 128

OpenStudy (anonymous):

1/128 or 0.0078125 or 0.78125%

OpenStudy (saifoo.khan):

Eliza the great.

OpenStudy (anonymous):

you are the bomb

OpenStudy (saifoo.khan):

are u? :O

OpenStudy (anonymous):

:D saifoo the greater

OpenStudy (saifoo.khan):

Haha!

OpenStudy (anonymous):

I will be coming back with another,. :)

jimthompson5910 (jim_thompson5910):

I don't know if it's that simple since there is the condition that "at least one is a head". Wouldn't that alter the probability?

OpenStudy (anonymous):

It wasn't right, but that's ok

OpenStudy (anonymous):

well if there were two heads then the answer would obviously be 100%

OpenStudy (anonymous):

What are the odds in favor of drawing a card lower than a 4 from an ordinary deck of cards?

jimthompson5910 (jim_thompson5910):

It might be easier to compute the probability of getting 1 tail...hmm

OpenStudy (anonymous):

what?

jimthompson5910 (jim_thompson5910):

well just thinking out loud really

OpenStudy (anonymous):

i think what the problem is saying is that if the coin is flipped and it lands on heads 1 time what is the probability that it will land on heads 7 more times but that would just be a distractor.

OpenStudy (anonymous):

Does anyone want to take a chance at this one?

OpenStudy (anonymous):

What are the odds in favor of drawing a card lower than a 4 from an ordinary deck of cards?

OpenStudy (anonymous):

with or without jokeers?

OpenStudy (anonymous):

that's all it says

jimthompson5910 (jim_thompson5910):

P(lower than 4) = 12/52 = 3/13 This is assuming that aces are considered lower than 2s. I'm also assuming no jokers are in the deck.

OpenStudy (anonymous):

well without jokers it would be 3x4 (because there are 3 numbers less than 3 and 4 suits) divided by 52

jimthompson5910 (jim_thompson5910):

So odds in favor are 3:10

OpenStudy (anonymous):

Thanks Jim...I do know the first answer is 2, but the homework is looking for 2 to _____

jimthompson5910 (jim_thompson5910):

oh so the number 2 is in the odds?

OpenStudy (anonymous):

Well, as far as I know probabilities are a HUGE fail for me... I love math, but this makes me hate it

OpenStudy (anonymous):

yah

jimthompson5910 (jim_thompson5910):

well then aces are not lower than 2s so P(lower than 4) = 8/52 = 2/13 so odds in favor are 2:11

OpenStudy (anonymous):

you are the BOMB!!!! Let me see, BRB

OpenStudy (anonymous):

you are correct.

jimthompson5910 (jim_thompson5910):

good to know, it would have been better if they let you know the details about the cards (eg: if jokers are included and if aces are lower than 2s, etc)

OpenStudy (anonymous):

ThanksJim...Are you up for a couple more?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Ok here's another for yo... Assume that the cost to spin the wheel once is $5.00 and that you will receive the amount shown on the spinner after it stops. The spinner has 3 separate section $2.00 is the biggest half, then $8.00, then $5.00. The spinner is stopped in the $2.00 half

OpenStudy (anonymous):

Let me know if that doesn't make sense

jimthompson5910 (jim_thompson5910):

how is the board divided up? Ie how much area does each sector have?

OpenStudy (anonymous):

one sec...

OpenStudy (anonymous):

The $2.00 half is half of the circle on left side. The $8.00 half is about 3/4 of the other half with the $5.00 section making up the rest

jimthompson5910 (jim_thompson5910):

So P($2) = 1/2, P($8) = (3/4)(1/2) = 3/8 and P($5) = 1-(1/2+3/8) = 1-7/8 = 1/8 What is the question asking again?

jimthompson5910 (jim_thompson5910):

ok I see it

OpenStudy (anonymous):

The website I gave you it's number 45

jimthompson5910 (jim_thompson5910):

I'm going to assume that the $8 portion is 1/3 of the board

jimthompson5910 (jim_thompson5910):

so the $5 portion is 1/2 - 1/3 = 1/6 of the board

jimthompson5910 (jim_thompson5910):

and the $2 region is 1/2 the board

OpenStudy (anonymous):

yah the 1/8 wasn't right :)

jimthompson5910 (jim_thompson5910):

Now let's compute expected value: Expected value = sum of probability*cost Expected value = 8(1/3) + 5(1/6) + 2(1/2) Expected value = 8/3 + 5/6 + 1 Expected value = 27/6 Expected value = 9/2 Expected value = 4.5 So you expect to get $4.50 on an average spin. But the cost is $5 per game, so you really lose 5-4.5 = 0.5 dollars or 50 cents per game. So the game is NOT fair. Note: a "fair game" is one where you neither gain nor lose money on average. Nobody wins and nobody loses.

OpenStudy (anonymous):

So E = $0.50

OpenStudy (anonymous):

It didn't come back as correct

jimthompson5910 (jim_thompson5910):

should be -0.50

jimthompson5910 (jim_thompson5910):

because you're losing 50 cents on average, what did the answer say?

OpenStudy (anonymous):

I put the - in front and that worked. Jim, um would you like to be my tutor for next weeks quiz? :) You are so wonderful!!

OpenStudy (anonymous):

These killed me this week and that's an understatement. I can't grasp the concept of it all

jimthompson5910 (jim_thompson5910):

Sure, I'd love to help you out.

OpenStudy (anonymous):

Do you want to give me your email address or no? Understand if not.

jimthompson5910 (jim_thompson5910):

let me know when you get it

OpenStudy (anonymous):

goti t

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

What are the expectations for the following $1 bets on a U.S. roulette wheel? 1 to 18 Two-number bet Five-number bet Double 00

OpenStudy (anonymous):

This is the last one tonight if you want to help me with this one

jimthompson5910 (jim_thompson5910):

I'm not familiar with roulette terminology or how the game is played. Where does it explain it in the book you showed me?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

Page 477 from the same website

jimthompson5910 (jim_thompson5910):

ok I see it

jimthompson5910 (jim_thompson5910):

ok I think I have the basics down, but it doesn't say anything about 1 to 18 or anything like that. I'm assuming those are special game types right?

OpenStudy (anonymous):

If you look at the example of the roulette table you will see 1 to 18 on the top right almost under 0 in that's ingreen

jimthompson5910 (jim_thompson5910):

Ok I think I found it. It says it pays 1 to 1. Is that what you're referring to?

OpenStudy (anonymous):

Haha! I don't know.

jimthompson5910 (jim_thompson5910):

Hmm sry I don't think I'm reading the right thing then.

OpenStudy (anonymous):

Well, don't worry about it. I am done for the night and will be trying again tomorrow. I will be in contact. Thanks so much for all you've done.

jimthompson5910 (jim_thompson5910):

ok glad to be of help, cya later